Consider the perfect competitive market for diesel.The aggregate demand for gasoline is Qd=100-p ,While the aggregate supply is Qs=3p

a)Work out the equilibrium price and quantity.

b) at the equilibrium level established in part a calculate the consumer surplus, producer surplus and total surplus.

c) it is established that the most fuel stations are going out of business to address this problem, the government decides to set a minimum price of p==30. What will be the new equilibrium price and quantity? What will be the new consumer surplus and producer surplus? Who gains and who loses from this regulation? Explain how the total surplus is affected? Briefly explain the intuition.


1
Expert's answer
2021-09-06T07:06:41-0400
"100-p=3p"

"4p=100"

"p=25"

"Q=75"

"\varDelta Q_D=\frac{100-25}{2}\times75=2812.5"

"\varDelta Q_S=\frac{25-0}{2}\times75=937.5"

If p=30


"\varDelta Q_D=\frac{100-30}{2}\times70=2450"

"\varDelta Q_S=\frac{100-30}{2}\times90=3150"

The increase in prices will be beneficial to the manufacturer.